Three fair dice, one white, one red and one black, are tossed and the number of spots on the top faces are noted. Which of the following have probability 1/4 :
(a) The red die is a six or the black die is even.
(b) All three numbers are even.
(c) The red die is even and the black die is 4 or more.
(d) Exactly two of the three numbers are equal.
(e) None of the above

Answer by stanbon(75887)

Three fair dice, one white, one red and one black, are tossed and the number of spots on the top faces are noted. Which of the following have probability 1/4 :
----
There are 6^3 = 216 possible outcomes
--------------------------
(a) The red die is a six or the black die is even.
Ways to succeed: 1*3*6 = 18
P(event) = 18/216 is not 1/4
----------------
(b) All three numbers are even.
Ways to succeed: 3*3*3 = 27
P(event) = 27/216 is not 1/4
------------------------------------
(c) The red die is even and the black die is 4 or more.
Ways to succeed: 3*3*6 = 54
P(event) = 54/216 = 1/4
-------------------------------------
(d) Exactly two of the three numbers are equal.
Ways to succeed: 6*6*5 = 180
P(event) = 180/216 is not 1/4
----------------------
(e) None of the above----no
==================================
cheers,
Stan H.
===============